OpenStudy (anonymous):

What is the equation of the line passing through (–3, 6) and parallel to the line y = 4x – 2 in standard form?

OpenStudy (anonymous):

need help

OpenStudy (bahrom7893):

y-6=4(x+3) y=4x+12+6 -4x+y=18

Directrix (directrix):

Definitions: Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers. 4x + (-1) y = -18 http://jwilson.coe.uga.edu/emt668/emat6680.2002/jackson/chapter%205%20lesson%20plan/day6.html

OpenStudy (bahrom7893):

uh oh thanks direct!

Directrix (directrix):

You did all the crinkum-crankum.

OpenStudy (bahrom7893):

i never had to deal with these stupid forms at my school, pheeww

Directrix (directrix):

I didn't either. There's a point of confusion about standard form for a quadratic and general form for a quadratic that has stymied me. I'll send over the link for you to view.